# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v2

@@ 3 Private static MySingleton mInstance;
Unreleased References. Another significant pattern can be characterized by the failure to release references (or have them garbage collected). For example, if a Fragment is held as an implicit reference, it won’t be garbage collected until the object holding the reference completes its execution. Initializing objects in complex Activity lifecycle methods can also result in unreleased references, ending up with memory leak. Listing 2 demonstrates a code snippet from Stack Overflow, where a singleton object holds an mContext reference (line 2). Here, mContext is a layer that stands behind its component (Activity, Service, etc.), and its lifecycle is much shorter than that of the singleton object. This prevents it from being garbage collected, leading to a memory leak.
@@
Redundant Objects. The creation of redundant objects is a common pattern that causes performance issues. Developers often mistakenly create new instances repeatedly instead of referring to existing ones, leading to endless memory consumption. Additionally, when two classes are mutually dependent and create instances of each other during initialization, it results in endless object creation. Moreover, creating new objects with the same functionality can also unnecessarily degrade performance. Listing 3 demonstrates a code snippet from Stack Overflow, where every time setTypeface (line 2) is called, it creates a new Typeface object (line 4, 6, 8). This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes.
@@ 5 FileOutputStream fOutputStream = new FileOutputStream(newFile);
Large-Scale Data. Large-scale data is also a culprit for performance issues. Loading large images without proper caching can cause excessive memory usage and slowdowns. Uploading large files without separating them into smaller chunks can lead to high memory consumption. As shown in a code snippet from Stack Overflow, where the buffer size is the same as the uploaded file. If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error.
@@
UI Operations. UI is another key factor that can affect performance, encompassing issues such as repeatedly drawing the same content and UI layout misconfiguration. For example, in the code snippet 5 from Stack Overflow, the ‘for’ loop in ‘onDraw’ (line 5) redraws the same circles every time when onDraw is invoked, even though the UI content doesn’t change. This causes the application to slow down and increases the computational load.
Other Patterns. Besides the previously mentioned patterns, other types of patterns also warrant attention, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager on the main thread, etc.
@@ 3. Studies on Performance Issues
In the above Section 2, we systematically examined real-world Android performance issues, their root causes, and common code patterns. However, it remains unclear whether cutting-edge literature targets the same practical issues, whether existing techniques can effectively resolve them, and whether current datasets can comprehensively evaluate the techniques’ effectiveness. To answer these questions, we conducted a systematic literature review (SLR) to assess the current state of Android performance analysis. By comparing the performance issues, techniques, and datasets discussed in research papers with those observed in real-world discussions, we aim to uncover gaps between academic research and real-world challenges.
@@ 3.1. Literature Review
To conduct a systematic literature review, we followed the methodologies outlined by Brereton et al.  (brereton2007lessons, ) and Kitchenham et al.  (Kitchenham2004ProceduresFP, ). Fig. 3 illustrates the working process which comprises 5 steps.
@@ 3.1.2. Repository Search
To ensure systematic literature collection, we conducted keyword-based searches across five libraries: IEEE Library, ACM Library, Springer Library, Science Direct Library, and Wiley Library. From this initial search, we retrieved a total of 11,442 papers. However, many of these studies were outside the scope of our focus. To narrow down to the most relevant papers, we concentrated on the fields of software engineering and mobile, selecting research codes 4606 and 4612 for conferences, and 0805 and 0803 for journals, specifically targeting publications containing the keywords “software” and “mobile”. In addition, we exclusively chose those published from conferences and journals rated CORE A*/A [^2] CORE provides conference and journal ratings. For conference: https://portal.core.edu.au/conf-ranks/. For journal: https://portal.core.edu.au/jnl-ranks/ from 2012 to 2024. We then systematically reviewed each of these venues, applying the aforementioned keywords to search for the titles of relevant publications. As a result, we successfully identified 441 publications distributed across 23 venues as listed in Table 2. Notably, there is no relevant paper identified in ECSA, FOSSACS, and SEAMS.
@@ 3.1.3. Paper Exclusion
To gather relevant papers as possible, we initially considered all returned results from our search. However, not all papers aligned with our focus on performance issues in mobile apps caused by poor coding practices, as discussed in Section 1. To refine our collection, we conducted a thorough review of the abstracts (and full content when necessary) of these papers, aiming to retain only those closely related to our research focus. Specifically, we applied the following exclusion criteria to streamline this process:
  1. (1)
    Short papers (i.e., less than six pages in double-column format or 11 pages in single-column format) were excluded.
@@
After applying these exclusion criteria, we retained 49 papers that are closely related to the topic of performance issues analysis at the Android mobile level.
@@ 3.1.4. Back/Forward Snowballing
After initially identifying relevant papers (i.e., 49 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature. The backward snowballing process involved scrutinizing the references of identified papers for additional relevant sources. Additionally, we performed the forward snowballing method by examining papers citing the identified papers. We strictly adhered to the exclusion criteria when conducting snowballing until no further papers could be identified. Eventually, we were able to locate a total of 66 papers.
@@ 3.1.5. Cross Check
Additionally, to avoid overlooking any primary studies, we conducted cross-checks involving at least two authors, ensuring the reliability of these papers and enhancing the trustworthiness of the primary studies. Ultimately, we collected 66 papers to assess the current state of research and identify gaps in addressing real-world performance issues.
@@ 3.2.1. Result - Targeted Performance Issues
Holistic Taxonomy. Through reviewing 66 research papers, we identified five types of performance issues and 44 contributing factors, marking each in the Lit. column of Table 1 with a checkmark (✓). By incorporating insights from both real-world discussions and research papers, we developed a final taxonomy encompassing seven types of performance issues and 72 factors. By examining these issues and their contributing factors from both real-world data and academic findings, we identified significant gaps between research focus and real-world challenges as follows.
Figure 4. Annual publications on each performance issue.
@@
Gaps between research focus and real-world performance issues. As shown in Fig. 4, we identified five key performance issues from literature: energy consumption, responsiveness, memory consumption, CPU usage, and GPU usage. These performance issues are also prevalent in real-world scenarios, indicating that researchers have made significant efforts to address concerns that are relevant to both developers and users. Despite these efforts, there remains a noticeable imbalance in focus. Among the five issues, energy consumption received the most attention, with 53 of the examined papers (80.30%) focusing on this issue, and this trend is likely to continue. However, despite this focus on energy consumption, Section 2.2 revealed that users are more concerned with responsiveness, while developers prioritize memory consumption. This highlights a discrepancy between the performance issues prioritized by researchers, developers, and users. Furthermore, two performance issues prevalent in real-world scenarios have not been explored in literature: storage consumption and internet data usage. These gaps suggest that researchers should place greater emphasis on the performance issues that are of primary concern to both users and developers in real-world apps.
@@
Gaps in academic research on real-world performance issue factors. By analyzing the cases where both the R-W. and Lit. columns are marked with a checkmark (✓), we found that only 21 out of 49 factors discussed in real-world settings (42.86%) have been studied in literature. It indicates that the majority of performance issues encountered in real-world scenarios remain underexplored in academic research. For instance, memory consumption, which is the top concern for developers, real-world data shows 20 contributing factors, but only 8 (40.00%) have been explored in literature. Similarly, for responsiveness, the top concern for users, 14 real-world factors were identified, but only nine (64.29%) have been studied in literature. Less common issues like CPU usage and GPU usage have seen even less attention, with only 25.00% and 50.00% of their respective real-world root causes being studied in literature. These gaps suggest that many root causes of real-world performance issues still remain unexplored by researchers. Notably, all identified root causes of energy consumption have been addressed in literature, indicating that this area has been extensively studied. This observation aligns with our finding that the majority of research efforts are concentrated on energy consumption.
@@
The presence of research-only factors not seen in real-world discussions. By examining rows where the R-W. column is unmarked (✗) but the Lit. column is marked (✓), we identified 23 factors that have been explored in literature but not observed in the real world. Specifically, we found one unseen factor in memory consumption, four in responsiveness, and 18 in energy consumption. This suggests that literature may uncover factors that did not appear in our real-world exploration, thereby further enriching the taxonomy. Notably, most of these unseen factors pertain to energy consumption, accounting for 90.00% of the factors explored in literature but not observed in the real-world. For example, Li et al. (Li2014MakingWA, ) found that OLED displays consume more energy when displaying light colors than dark colors. Similarly, Cito et al. (Cito2016BatteryawareTI, ) observed that frequent advertisement and analytics (A&A) service requests lead to significant energy waste. On the other hand, there is a substantial number of energy consumption factors that appear only in the literature, yet rarely manifest in real-world applications. This indicates a pronounced disconnection between academic research and the practical challenges faced by practitioners.
@@
Answers to RQ2.1: Through an extensive integration of real-world explorations and thorough literature review, we present a comprehensive taxonomy encapsulating seven primary types of performance issues and 72 contributing factors. A noteworthy finding shows that only 42.86% factors contributing to performance issues have been examined in academic research. A significant misalignment exists in the primary concerns of researchers, developers, and users: researchers concentrate on energy consumption, developers prioritize memory usage, and users are chiefly concerned with responsiveness.
@@ 3.2.2. Result - Technique
To address performance issues, researchers have proposed automatic tools using diverse techniques, categorized into static approaches, dynamic approaches, and deep learning approaches. Table 3 presents the list of available automatic approaches collected from our examined papers. Among the 66 papers examined, eight (12.12%) provided publicly accessible tools designed to detect specific factors contributing to particular performance issues. This small fraction can be attributed to several reasons: some papers lack accessible artifacts, others provide only datasets without the corresponding tool code, and in some cases, the links to the artifacts are invalid. This indicates a significant gap between automatic tools with their practical usage to within our community.
In addition, we found that existing tools target only 17 out of the 72 factors in our taxonomy (Table 1), covering just 23.61%. This indicates that the majority of factors contributing to the seven primary types of performance issues are overlooked by SOTA tools, highlighting a significant gap in addressing these issues comprehensively. Moreover, most automated tools are primarily designed to detect performance issues rather than actively resolve them. Given that these detection tools rely on specific code patterns, we can anticipate that the identified patterns can also be applied to remediate performance issues. We then summarize the approaches as follows:
@@
Tool Approach Technique Stage Performance Issue Target Factor Link
VALA (Lu2022DetectingRU, ) Static Dataflow Analysis Symbolic Execution Detect Memory Consumption Energy Consumption Activity Resource  (Lu2022DetectingRUTool, )
AsyncChecker(Pan2020StaticAC, ) Static Pattern Checking Interprocedural- Code Analysis Detect Memory Consumption Energy Consumption Responsiveness AsyncTask  (Pan2020StaticACTool, )
TAPIR (Li2019CharacterizingAD, ) Static Code Pattern Analysis Detect Memory Consumption Energy Consumption Responsiveness Image  (Li2019CharacterizingADTool, )
ServDroid (Song2019ServDroidDS, ) Static Code Pattern Analysis Interprocedural Code Analysis Detect Memory Consumption Energy Consumption Service  (Song2019ServDroidDSTool, )
Lyu et al. (Lyu2018RemoveRF, ) Static Dataflow Analysis Program transformation Detect Fix Energy Consumption Responsiveness Database Operation Loop  (Lyu2018RemoveRFTool, )
PlumbDroid (Bhatt2020AutomatedRO, ) Static Control-Flow Analysis Patch Generation Detect Fix Responsiveness Resource  (Bhatt2020AutomatedROTool, )
µDroid (Jabbarvand2017DroidAE, ) Dynamic Mutation Testing Power Trace Analysis Detect Energy Consumption Resource Wakelock Loop  (Jabbarvand2017DroidAETool, )
ACETON (Jabbarvand2020AutomatedCO, ) DL Dynamic Deep Learning Feature Engineering Runtime Monitoring Detect Energy Consumption Resource  (Jabbarvand2020AutomatedCOTool, )
@@
Static Analysis. Most existing approaches are based on static analysis to detect or fix performance issues, as shown in Table 3. This is intuitive since static analysis involves reasoning about program behavior without execution, making it lightweight to deploy Specifically, taint analysis, symbolic execution (Lu2022DetectingRU, ), code pattern analysis (Pan2020StaticAC, ; Li2019CharacterizingAD, ; Song2019ServDroidDS, ), and data flow analysis (Lyu2018RemoveRF, ) are commonly employed in rule-based static analysis as many performance issues arise from common code patterns. Additionally, static approaches, such as feature engineering, are used to assist deep learning approaches by extracting semantic code features (Kim2019AMD, ; Su2016ADL, ). Moreover, they can also provide fundamental information for dynamic fuzzing tools to trigger performance bugs on the fly (Guo2022DetectingAF, ). Furthermore, static approaches play an important role in fixing performance issues, particularly in generating patches for automatic program repair (Lyu2018RemoveRF, ; Bhatt2020AutomatedRO, ).
@@
Dynamic Analysis. There are only two studies that primarily employ dynamic analysis approaches to detect performance bugs. This result aligns with expectations because dynamic analysis involves executing apps in a controlled environment and is unable to cover all possible paths (dong2020timemachine, ). However, dynamic analysis is more resilient to evasion techniques (e.g., reflection (sun2021taming, ), obfuscation (sun2023demystifying, )) and can supplement static approaches to eliminate false positives (sun2022mining, ). As shown in Table 3, both dynamic analysis tools target energy consumption issues, particularly unnecessary energy waste such as resource under-utilization, wakelock misuse, and inefficient loops, which are detectable through abnormal energy consumption monitoring. We also discovered one study (Jabbarvand2020AutomatedCO, ) that employed dynamic monitoring to collect features for deep learning methodologies aimed at detecting energy issues caused by resource inefficiencies.
@@
Answers to RQ2.2: We identified eight publicly available tools, primarily using static analysis, dynamic analysis, and deep learning approaches. Existing tools target only 17 of the 72 identified factors to performance issues (23.61%), leaving a substantial number of factors unexamined. Most tools focus on detecting performance issues while neglecting their resolution.

### v3

@@ 3 Private static MySingleton mInstance;
Unreleased References. Another significant pattern can be characterized by the failure to release references (or have them garbage collected). For example, if a Fragment is held as an implicit reference, it won’t be garbage collected until the object holding the reference completes its execution. Initializing objects in complex Activity lifecycle methods can also result in unreleased references, ending up with memory leak. Listing 2 demonstrates a code snippet from Stack Overflow, where a singleton object holds an mContext reference (line 2). Here, mContext is a layer that stands behind its component (Activity, Service, etc.), and its lifecycle is much shorter than that of the singleton object. This prevents it from being garbage collected, leading to a memory leak. This case mainly involves the Singleton, Activity, and Service factors.
@@
Redundant Objects. The creation of redundant objects is a common pattern that causes performance issues. Developers often mistakenly create new instances repeatedly instead of referring to existing ones, leading to endless memory consumption. Additionally, when two classes are mutually dependent and create instances of each other during initialization, it results in endless object creation. Moreover, creating new objects with the same functionality can also unnecessarily degrade performance. Listing 3 demonstrates a code snippet from Stack Overflow, where every time setTypeface (line 2) is called, it creates a new Typeface object (line 4, 6, 8). This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes. This case mainly involves the Object Creation factor.
@@ 5 FileOutputStream fOutputStream = new FileOutputStream(newFile);
Large-Scale Data. Large-scale data is also a common cause of performance issues, since operations on such data often exceed the capacity of mobile devices with limited resources. Loading large images without proper caching can cause excessive memory usage and slowdowns. Uploading large files without separating them into smaller chunks can lead to high memory consumption. As shown in a code snippet from Stack Overflow, where the buffer size is the same as the uploaded file. If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error. This case mainly involves the Massive Data and File Handling factors.
@@
UI Operations. UI is another key factor that can affect performance, encompassing issues such as repeatedly drawing the same content and UI layout misconfiguration. For example, in the code snippet 5 from Stack Overflow, the ‘for’ loop in ‘onDraw’ (line 5) redraws the same circles every time when onDraw is invoked, even though the UI content doesn’t change. This causes the application to slow down and increases the computational load. This case mainly involves contributing factors such as Loop and UI Rendering.
Other Patterns. Besides the previously mentioned patterns, other types of patterns also warrant attention, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager on the main thread, etc. We made the full list publicly available in our artifact: (our_repo, ).
@@
2.3. Practitioner Validation of the Taxonomy
To further validate whether the performance issues and contributing factors identified in our taxonomy reflect practical concerns, we conducted a survey with Android practitioners.
2.3.1. Method
We designed a questionnaire consisting of the following parts:
  • Q1a: What is your current role related to Android software development?
  • Q1b: What is your age range? (23 or below, 24–27, 28 or above)
  • Q1c: How many years of experience do you have in Android development or related research?
  • Q2: This section presents seven common categories of performance issues. For each category, please indicate the extent to which it concerns you during development. Responses were recorded on a 5-point Likert scale: 1 = not concerned at all, 2 = slightly concerned, 3 = moderately concerned, 4 = concerned, 5 = very concerned.
  • Q3: This section presents each performance issue category along with its contributing factors and an example illustrating how each factor can lead to the issue. For each category, you are asked to review the associated factors and select all those that they consider a concern or believe may potentially affect performance during development. This question used a multiple-choice format, allowing participants to select zero or more factors that reflect their concerns or are potentially relevant to their development experience.
Among them, Q2 was designed to validate the relevance of the seven performance issue categories by measuring how strongly participants with Android development experience perceive each category as a practical concern. Q3 was designed to validate the comprehensiveness of the contributing factors by examining whether participants recognize these factors as relevant to their own development experience, thereby assessing the alignment between our taxonomy and real-world practice. The survey was conducted online and remained open for two weeks. We collected anonymous responses to encourage honest feedback, and all participants provided informed consent before answering the questions.
Figure 3. Average concern scores for different categories of Android performance issues.
2.3.2. Results
We received 21 valid responses. Participants’ ages fell into three ranges: 23 or below (4.8%), 24–27 (76.2%), and 28 or above (19.0%), with the majority in the 24–27 range. Most participants (81%) are casual developers with Android development experience, while the remaining 19% are professional developers. On average, participants reported 4.19 years of experience in Android development or related research. For Q2, the median concern score across all seven performance issue categories was 4, indicating that participants widely recognize these issues as important in practice. Fig. 3 shows the average concern level for each performance issue category. The most concerning issue was Memory Consumption (mean: 4.95), while the least concerning was Internet Data Usage, which aligns with the patterns we observed in real-world discussions. For Q3, participants selected an average of 54.19 out of 63 contributing factors, resulting in an overall alignment rate of 86.16%. Notably, 18 factors were selected by all participants, such as the UI Rendering factor contributing to GPU Usage issues and the Resource factor contributing to Energy Consumption problems. This suggests strong agreement between our taxonomy and developers’ real-world experiences. In contrast, some factors received fewer selections. For example, the Third-party Library factor contributing to CPU Usage issues and the Singleton factor contributing to Memory Consumption issues were each selected by only 11 participants. This may be because these factors occur less frequently or are less visible during development, leading to them being unintentionally overlooked by developers.
Overall, the survey results support the validity of our taxonomy and confirm that the performance issues and contributing factors identified in our study accurately reflect real-world development concerns, thereby reinforcing the findings presented in RQ1.
@@ 3. Studies on Performance Issues
In the above Section 2, we systematically examined real-world Android performance issues, their root causes, and common code patterns. However, it remains unclear whether cutting-edge literature targets the same practical issues (RQ2.1), whether existing techniques can effectively resolve them (RQ2.2), and whether current datasets can comprehensively evaluate the techniques’ effectiveness (RQ2.3). To answer these questions, we conducted a systematic literature review (SLR) to assess the current state of Android performance analysis. By comparing the performance issues, techniques, and datasets discussed in research papers with those observed in real-world discussions, we aim to uncover gaps between academic research and real-world challenges.
@@ 3.1. Literature Review
To conduct a systematic literature review, we followed the methodologies outlined by Brereton et al.  (brereton2007lessons, ) and Kitchenham et al.  (Kitchenham2004ProceduresFP, ). Fig. 4 illustrates the working process which comprises 5 steps. Here, we clarify that this study focuses on performance issues resulting from inefficient coding practices within Android applications, rather than problems related to Android device hardware or the operating system.
@@ 3.1.2. Repository Search
To ensure systematic literature collection, we conducted keyword-based searches across five libraries: IEEE Library, ACM Library, Springer Library, Science Direct Library, and Wiley Library. From this initial search, we retrieved a total of 6,649 papers from IEEE, 1,507 from ACM, 537 from Springer, 1,969 from ScienceDirect, and 1,008 from Wiley. Since many papers appeared across multiple libraries, we removed duplicates based on paper titles, resulting in 11,574 unique papers. However, many of these studies were outside the scope of our focus. To narrow down to the most relevant papers, we concentrated on the fields of software engineering and mobile, selecting research codes 4606 and 4612 for conferences, and 0805 and 0803 for journals, specifically targeting publications containing the keywords “software” and “mobile”. In addition, we exclusively chose those published from conferences and journals rated CORE A*/A [^2] CORE provides conference and journal ratings. For conference: https://portal.core.edu.au/conf-ranks/. For journal: https://portal.core.edu.au/jnl-ranks/ from 2012 to 2024. We then systematically reviewed each of these venues, applying the aforementioned keywords to search for the titles of relevant publications. As a result, we successfully identified 441 publications distributed across 23 venues as listed in Table 2. Notably, there is no relevant paper identified in ECSA, FOSSACS, and SEAMS.
@@ 3.1.3. Paper Exclusion
To gather as many relevant papers as possible, we initially considered all returned results from our search. However, not all papers aligned with our focus on performance issues in mobile apps caused by poor coding practices, as discussed in Section 1. To refine our collection, we conducted a thorough review of the abstracts (and full content when necessary) of these papers, aiming to retain only those closely related to our research focus. Specifically, we applied the following exclusion criteria to streamline this process:
@@
After applying these exclusion criteria, we retained 60 papers that are closely related to the topic of performance issues analysis at the Android mobile level.
@@ 3.1.4. Back/Forward Snowballing
After initially identifying relevant papers (i.e., 60 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature. The backward snowballing process involved scrutinizing the references of identified papers for additional relevant sources. Additionally, we performed the forward snowballing method by examining papers citing the identified papers. We strictly adhered to the exclusion criteria when conducting snowballing until no further papers could be identified. Eventually, we were able to locate a total of 85 papers.
@@ 3.1.5. Cross Check
Additionally, to avoid overlooking any primary studies, we conducted cross-checks involving at least two authors, ensuring the reliability of these papers and enhancing the trustworthiness of the primary studies. Ultimately, we collected 85 papers to assess the current state of research and identify gaps in addressing real-world performance issues. As shown in Fig. 5(a), the majority of these studies were published in conference proceedings (62 out of 85, accounting for 72.94%). Moreover, we observe a consistent research interest in this topic, with papers published every year since 2012. All papers involved in the SLR process are available in our replication package under the Literature_Review/data directory (our_repo, ).
@@ 3.2.1. Result - Targeted Performance Issues
Holistic Taxonomy. Through reviewing 85 research papers, we identified five types of performance issues and 45 contributing factors, marking each in the Literature column of Table 1 with a checkmark (✓). By incorporating insights from both real-world discussions and research papers, we developed a final taxonomy encompassing seven types of performance issues and 82 factors. By examining these issues and their contributing factors from both real-world data and academic findings, we identified significant gaps between research focus and real-world challenges as follows.
(a) Number of publications per year by venue type.
(b) Annual publications on each performance issue.
Figure 5. Distribution of selected studies by venue type and performance issue.
@@
Gaps between research focus and real-world performance issues. As shown in Fig. 5(b), we identified five key performance issues from literature: energy consumption, responsiveness, memory consumption, GPU usage, and CPU usage. These performance issues are also prevalent in real-world scenarios, indicating that researchers have made significant efforts to address concerns that are relevant to both developers and users. Despite these efforts, there remains a noticeable imbalance in focus. Among the five issues, energy consumption received the most attention, with 69 of the examined papers (81.18%) focusing on this issue, and this trend is likely to continue. However, despite this focus on energy consumption, Section 2.2 revealed that users are more concerned with responsiveness, while developers prioritize memory consumption. This highlights a discrepancy between the performance issues prioritized by researchers, developers, and users. Moreover, two performance issues that are prevalent in real-world applications, storage consumption and internet data usage, have been relatively underexplored in the academic literature (Bijlani2021WhereDM, ; Zhang2012HowEA, ; Liu2016WhoMM, ), highlighting the need for more focused attention in future research. These gaps suggest that researchers should place greater emphasis on the performance issues that are of primary concern to both users and developers in real-world apps.
@@
Gaps in academic research on real-world performance issue factors. By analyzing the cases where both the Real-World and Literature columns are marked with a checkmark (✓), we found that only 27 out of 63 factors discussed in real-world settings (42.86%) have been studied in literature. It indicates that the majority of performance issues encountered in real-world scenarios remain underexplored in academic research. For instance, memory consumption, which is the top concern for developers, real-world data shows 21 contributing factors, but only 8 (38.10%) have been explored in literature. Similarly, for responsiveness, the top concern for users, 16 real-world factors were identified, but only 10 (62.50%) have been studied in literature. Less common issues like CPU usage and GPU usage have seen even less attention, with only 14.29% and 50.00% of their respective real-world root causes being studied in literature. These gaps suggest that many root causes of real-world performance issues still remain unexplored by researchers. Notably, all identified root causes of energy consumption have been addressed in literature, indicating that this area has been extensively studied. This observation aligns with our finding that the majority of research efforts are concentrated on energy consumption.
@@
The presence of research-only factors not seen in real-world discussions. By examining rows where the Real-World column is unmarked (✗) but the Literature column is marked (✓), we identified 19 factors that have been explored in literature but not observed in the real world. Specifically, we found one unseen factor in memory consumption, four in responsiveness, and 14 in energy consumption. This suggests that literature may uncover factors that did not appear in our real-world exploration, thereby further enriching the taxonomy. Notably, most of these unseen factors pertain to energy consumption, accounting for 73.68% of the factors explored in literature but not observed in the real-world. For example, Li et al. (Li2014MakingWA, ) found that OLED displays consume more energy when displaying light colors than dark colors. Similarly, Cito et al. (Cito2016BatteryawareTI, ) observed that frequent advertisement and analytics (A&A) service requests lead to significant energy waste. On the other hand, there is a substantial number of energy consumption factors that appear only in the literature, yet rarely manifest in real-world applications. This indicates a pronounced disconnection between academic research and the practical challenges faced by practitioners.
@@
Answers to RQ2.1: Through an extensive integration of real-world explorations and thorough literature review, we present a comprehensive taxonomy encapsulating seven primary types of performance issues and 82 contributing factors. A noteworthy finding shows that only 42.86% of factors contributing to performance issues have been examined in academic research. A significant misalignment exists in the primary concerns of researchers, developers, and users: researchers concentrate on energy consumption, developers prioritize memory usage, and users are chiefly concerned with responsiveness.
@@ 3.2.2. Result - Technique
To address performance issues, researchers have proposed automatic tools using diverse techniques, categorized into static approaches, dynamic approaches, and deep learning approaches. Table 3 presents the list of available automatic approaches collected from our examined papers. Among the 85 papers examined, 14 (16.47%) provided publicly accessible tools designed to detect specific factors contributing to particular performance issues. This small fraction can be attributed to several reasons: some papers lack accessible artifacts, others provide only datasets without the corresponding tool code, and in some cases, the links to the artifacts are invalid. This indicates a significant gap between automatic tools and their practical usability within the community.
In addition, we found that existing tools target only 30 out of the 82 factors in our taxonomy (Table 1), covering just 36.59%. This indicates that the majority of factors contributing to the seven primary types of performance issues are overlooked by SOTA tools, highlighting a significant gap in addressing these issues comprehensively. Moreover, most automated tools are primarily designed to detect performance issues rather than actively resolve them. Given that these detection tools rely on specific code patterns, we can anticipate that the identified patterns can also be applied to remediate performance issues. We then summarize the approaches as follows:
@@
Tool Approach Technique Stage
Performance Issue
Target Factor
Link
VALA (Lu2022DetectingRU, ) Static
Dataflow Analysis
Symbolic Execution
Detect
Memory Consumption
Energy Consumption
Activity
Resource
 (Lu2022DetectingRUTool, )
AsyncChecker(Pan2020StaticAC, ) Static
Code Pattern Analysis
Interprocedural-
Code Analysis
Detect
Memory Consumption
Energy Consumption
Responsiveness
AsyncTask  (Pan2020StaticACTool, )
TAPIR (Li2019CharacterizingAD, ) Static
Code Pattern Analysis
Detect
Memory Consumption
Energy Consumption
Responsiveness
Image  (Li2019CharacterizingADTool, )
ServDroid (Song2019ServDroidDS, ) Static
Code Pattern Analysis
Interprocedural Code Analysis
Detect
Memory Consumption
Energy Consumption
Service  (Song2019ServDroidDSTool, )
Lyu et al. (Lyu2018RemoveRF, ) Static
Dataflow Analysis
Program Transformation
Detect
Fix
Energy Consumption
Responsiveness
Database Operation
Loop
 (Lyu2018RemoveRFTool, )
PlumbDroid (Bhatt2020AutomatedRO, ) Static
Control-Flow Analysis
Patch Generation
Detect
Fix
Responsiveness Resource  (Bhatt2020AutomatedROTool, )
µDroid (Jabbarvand2017DroidAE, ) Dynamic
Mutation Testing
Power Trace Analysis
Detect Energy Consumption
Resource
Wakelock
Loop
 (Jabbarvand2017DroidAETool, )
ACETON (Jabbarvand2020AutomatedCO, )
DL
Dynamic
Deep Learning
Feature Engineering
Runtime Monitoring
Detect Energy Consumption Resource  (Jabbarvand2020AutomatedCOTool, )
Leafactor (Cruz2017LeafactorIE, ) Static
Code Pattern Analysis
Program Transformation
Detect
Fix
Energy Consumption
Wakelock
View
UI Rendering
 (LeafactorTool, )
PAPRIKA (Hecht2015DetectingAI, ) Static
Code Pattern Analysis
Detect
Memory Consumption
Responsiveness
Energy Consumption
Inefficient Method
Data Type
Activity
 (PAPRIKATool, )
ecoCode (LeGoaer2022ecoCodeAS, ) Static
Rule-Based Detection
XML/Gradle Inspection
Detect Energy Consumption
Resource
Wakelock
Redundant Frames
Service
Loop
HTTP Request
 (ecoCodeTool, )
aDoctor (Iannone2020RefactoringAE, )
Static
Code Pattern Analysis
Program Transformation
Detect
Fix
Energy Consumption
Wakelock
Inefficient Method
Data Type
 (adoctorTool, )
Das et al. (Das2020CharacterizingTE, )
Static
Rule-Based Detection
AST Analysis
Detect
Energy Consumption
Memory Consumption
Responsiveness
UI Rendering
Data Type
Main Thread
Inefficient Method
View
Wakelock
Activity
 (AndroidLintTool, )
Vásquez et al. (Vsquez2015HowDD, )
Dynamic
Runtime Monitoring
Detect
Responsiveness
Main Thread
 (StrictModeTool, )
@@
Static Analysis. Most existing approaches are based on static analysis to detect or fix performance issues, as shown in Table 3. This is intuitive since static analysis involves reasoning about program behavior without execution, making it lightweight to deploy. Specifically, taint analysis, symbolic execution (Lu2022DetectingRU, ), code pattern analysis (Pan2020StaticAC, ; Li2019CharacterizingAD, ; Song2019ServDroidDS, ), and data flow analysis (Lyu2018RemoveRF, ) are commonly employed in rule-based static analysis as many performance issues arise from common code patterns. Additionally, static approaches, such as feature engineering, are used to assist deep learning approaches by extracting semantic code features (Kim2019AMD, ; Su2016ADL, ). Moreover, they can also provide fundamental information for dynamic fuzzing tools to trigger performance bugs on the fly (Guo2022DetectingAF, ). Furthermore, static approaches play an important role in fixing performance issues, particularly in generating patches for automatic program repair (Lyu2018RemoveRF, ; Bhatt2020AutomatedRO, ).
@@
Dynamic Analysis. There are only three studies that primarily employ dynamic analysis approaches to detect performance bugs. This result aligns with expectations because dynamic analysis involves executing apps in a controlled environment and is unable to cover all possible paths (dong2020timemachine, ). However, dynamic analysis is more resilient to evasion techniques (e.g., reflection (sun2021taming, ), obfuscation (sun2023demystifying, )) and can supplement static approaches to eliminate false positives (sun2022mining, ). As shown in Table 3, two dynamic analysis tools target energy consumption issues, particularly unnecessary energy waste such as resource under-utilization, wakelock misuse, and inefficient loops, which are detectable through abnormal energy consumption monitoring. One dynamic analysis tool, StrictMode (StrictModeTool, ), targets responsiveness issues, particularly those caused by executing heavy operations on the main thread, which can lead to UI freezes and slow interactions. We also discovered one study (Jabbarvand2020AutomatedCO, ) that employed dynamic monitoring to collect features for deep learning methodologies aimed at detecting energy issues caused by resource inefficiencies.
@@
Answers to RQ2.2: We identified 14 publicly available tools, primarily using static analysis, dynamic analysis, and deep learning approaches. Existing tools target only 30 of the 82 identified factors to performance issues (36.59%), leaving a substantial number of factors unexamined. Most tools focus on detecting performance issues while neglecting their resolution.